Reduce the homogeneous equation to the separated one

(x+y)dx+(y-x)dy=0


1
Expert's answer
2022-01-09T16:38:12-0500
"(x+y)+(y-x)y'=0"

Let "y(x)=xu(x)." Then


"y'=u+xu'"

"x+xu+(ux-x)(u+xu')=0"

"x(1+u+(u-1)(u+xu')=0"

"1+u+u^2-u+xuu'-xu'=0"

"x(1-u)u'=1+u^2"

"\dfrac{1-u}{1+u^2}du=\dfrac{dx}{x}"

Integrate


"\int\dfrac{1-u}{1+u^2}du=\int \dfrac{dx}{x}"

"\tan^{-1}(u)-\dfrac{1}{2}\ln(1+u^2)=\ln x+C"

"\tan^{-1}(\dfrac{y}{x})-\dfrac{1}{2}\ln(1+\dfrac{y^2}{x^2})=\ln x+C"